OpenStudy (anonymous):

Please help Q5, Equation of all straight lines with algebraic sum of the intercepts fixed as k..

ganeshie8 (ganeshie8):

lets say y intercept = c so x intercept becomes : k-c because their sum must equal to k. got it so far ?

OpenStudy (anonymous):

so, y-(k-c)=mx+c? is this correct?

ganeshie8 (ganeshie8):

not exactly..

ganeshie8 (ganeshie8):

lets draw a picture

OpenStudy (anonymous):

um, how is it?

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

u see, it goes through two points (0,c) and (k-c, 0) now use the equation for a line going through two points (x1, y1) , (x2, y2) \[ y-y1 = \frac{y2-y1}{x2-x1}(x-x1)\] \[ y-c = \frac{0-c}{k-c-0}(x-0)\] \[ y-c = \frac{c}{c-k}(x)\]

ganeshie8 (ganeshie8):

you can keep the equation as it is. it should be ok. or else you can try simplifying further..